OpenStudy (anonymous):

help ? *Advance Algebra With Financial Applications * Miguel needs $34,100 to purchase a boat. How much money will he need to invest today in a savings account earning 4.1% interest, compounding quarterly, to have enough money to purchase the boat in 16 years? $2,605.60 $17,754.46 $28,948.17 $31,935.91

OpenStudy (anonymous):

I used this formula FV = P((1+r/n)^(n*t) - 1 )/( r/n ) but it's not giving me the right answer.

OpenStudy (cruffo):

that looks like the savings plan formula (the one you use when making monthly payments). I think you just need the compound interest formula

OpenStudy (anonymous):

Do u have considered time ? Here it is quarterly . Morecoer you shoud use compound intrest formula

OpenStudy (cruffo):

\[A = P\left(1+\frac{r}{n}\right)^{nt}\]

OpenStudy (amistre64):

\[B_n=B_ok^{4*t}-P\frac{1-k^{4*t}}{1-k}\] \[B_{16}=B_ok^{4*16}-0\frac{1-k^{4*16}}{1-k}\] \[34100=B_ok^{4*16}\] \[\frac{34100}{k^{4*16}}=B_o~:~k=1+.041/4\]

OpenStudy (cruffo):

Ah... $34,100 is the desired future value

OpenStudy (cruffo):

\[\large 34,100 = P\left(1+\frac{0.041}{4}\right)^{(4\cdot 16)}\] You need to solve for P because you need to find out how much he should invest right now to get $34,100 in 16 years.

OpenStudy (amistre64):

I tend to use B for balances, and P for Payments ... as opposed to A for Amount and P for Principal :)

OpenStudy (cruffo):

that would be wise, but most texts use A and P for the compound interest formula, or possible A = FV and P = PV

OpenStudy (cruffo):

Note that that answer is rounded to the nearest 10 cents!

OpenStudy (cruffo):

I got $17, 754.46 rounding to the nearest cent.

OpenStudy (anonymous):

Me too

OpenStudy (cruffo):

Ya!